15 1)  Thomson Co. produces and distributes semiconductors for use by computer manufacturers. Thomson Co. issued $1,200,000...

15

1)  Thomson Co. produces and distributes semiconductors for use by computer manufacturers. Thomson Co. issued $1,200,000 of 10-year, 12% bonds on May 1 of the current year at face value, with interest payable on May 1 and November 1. The fiscal year of the company is the calendar year.

Journalize the entries to record the following selected transactions for the current year. Refer to the Chart of Accounts for exact wording of account titles.

May 1 Issued the bonds for cash at their face amount.
Nov. 1 Paid the interest on the bonds.
Dec. 31 Recorded accrued interest for two months.

Journalize the transactions. Refer to the Chart of Accounts for exact wording of account titles.

PAGE 10

JOURNAL

ACCOUNTING EQUATION

DATE DESCRIPTION POST. REF. DEBIT CREDIT ASSETS LIABILITIES EQUITY

1

2

3

4

5

Adjusting Entries

6

7

2)

Entries for Issuing Bonds and Amortizing Discount by Straight-Line Method

On the first day of its fiscal year, Chin Company issued $18,300,000 of five-year, 6% bonds to finance its operations of producing and selling home improvement products. Interest is payable semiannually. The bonds were issued at a market (effective) interest rate of 7%, resulting in Chin Company receiving cash of $17,539,055.

a. Journalize the entries to record the following:

  1. Issuance of the bonds.
  2. First semiannual interest payment. The bond discount amortization, using the straight-line method, is combined with the semiannual interest payment. (Round your answer to the nearest dollar.)
  3. Second semiannual interest payment. The bond discount amortization, using the straight-line method, is combined with the semiannual interest payment. (Round your answer to the nearest dollar.)

For a compound transaction, if an amount box does not require an entry, leave it blank. Round your answers to the nearest dollar.

1. Cash
Discount on Bonds Payable
Bonds Payable
2. Interest Expense
Discount on Bonds Payable
Cash
3. Interest Expense
Discount on Bonds Payable
Cash
Answer #1
Question-1
Date Description Post Ref Debit Credit Assets Liabilities Equity
May 1 Cash 1200000
     Bonds payable 1200000
Nov 1 Interest expense 72000 =1200000*12%*6/12
        Cash 72000
Dec 31 Interest expense 24000 =1200000*12%*2/12
     Interest payable 24000
Question-2
1 Cash 17539055
Discount on Bonds Payable 760945
      Bonds Payable 18300000
2 Interest Expense 625095
     Discount on Bonds Payable 76095 =760945/10
      Cash 549000 =18300000*6%*6/12
3 Interest Expense 625095
     Discount on Bonds Payable 76095 =760945/10
      Cash 549000 =18300000*6%*6/12